unicorn35:

Complete the table below to solve the equation 2.5x − 10.5 = 64(0.5x). x f(x) = 2.5x − 10.5 g(x) = 64(0.5x) 2 3 4 5 6 A newspaper started an online version of its paper 14 years ago. In a recent presentation to stockholders, the lead marketing executive states that the revenues for online ads have more than doubled that of the revenues for printed ads since starting the online version of the paper. Use the graph below to justify the lead executive’s statement. Determine the approximate year that the two ad revenues were equal. See Text Version Two ocean beaches are being affected by erosion. The table shows the width, in feet, of each beach measured at high tide where 1995 is represented by year 0: Year number Western Beach width (in feet) Dunes Beach width (in feet) 0 100 20 5 90 45 10 80 70 11 78 75 12 76 80 15 70 95 Describe the patterns shown by the erosion data measurements shown for each of the beaches in the table. Between which years will the beaches have approximately the same width? Assuming these rates remain constant, what can you do to get a better approximation of when the two beaches will have the same width?

unicorn35:

i kind of already did the work to the equations i just need the answer for Use the graph below to justify the lead executive’s statement. Determine the approximate year that the two ad revenues were equal.

Jaynater:

If you want, you can graph these two functions and see where they may intersect in relation to your overall equation. Make sure to label correctly if you do that.

unicorn35:

unicorn35:

1 attachment
unicorn35:

1 attachment
Jaynater:

Using the "middle half.PNG", you can see where the two lines intersect. Estimate that value and that will be your answer for when the two were equal

Jaynater:

Remember to use your x-values for this estimate, as that's the amount of time.

unicorn35:

my teacher said the same thing at this point im lost

unicorn35:

2.5x-10.5=6 Multiply each side by ten 2.5x. 10 -10.5. 10 = 6. 10 Solve 25x- 105 = 60 Add 105 to both sides 25x- 105+ 105 = 60 +105 Simplified 25x=165 Divide by 25 on both sides 25x/25 = 165/25 X=6.6 In the graph during year 14, revenue from online ads is 4.2 million, while revenue from printed ads is 1.7 million. Which means the revenue from online ads have more than doubled revenue from printed ads at least by year 14. Both curves intersect between the seventh and eighth year, I think that in about year seven the two revenues will be equal, and both online ads and printed ads generated a number revenue of 2.2 million each. The Width of beach at 8pm = 36m Width of beach at 3am = (36-8) m =28 m The Average change in width A= change in width which mean the Width of beach at 8 pm -Width of beach at 3 a.m.0/ time in hours A= (36-28)/7 A= 8/7 m/hr A = 1.142 m /hr

Jaynater:

So, you see that the two lines cross over in the "middle half.PNG" right?

unicorn35:

thats literally all my work

unicorn35:

yes

Jaynater:

I know, I know. But we don't need that much to solve this, it's very simple. You just move over to the x value where they intersect (between the labeled 6 and 8 value, more closer to between 7 and 8) and that's your answer.

Jaynater:

So, again, approximately what year?

unicorn35:

year 14?

Jaynater:

No, not quite as that's not the X-Value (The little axis at the bottom) that our lines intersect/cross over at.

unicorn35:

ok so i know that by year seven the revenue will be equal what am i missing her specific words were You did not substitute the x values in the equation to find the solution for part A but was able to justify's the executive statement in part B. In addition, you did not answer the last two questions in part C. . . 

Jaynater:

Okay, I'm a little confused. I'm helping you with number 2) Part B. Are we on the same page?

unicorn35:

all three of those images are one assignment part one i answered showing my work up at the top then also at the top i answered part b but not part a i do not know the answer to part a so she said i did not answer that then i for got to substitute for ex and answer the questions down at the bottom for part c i just need the answer for those two and ill be done

Jaynater:

Well, if you can find a point in your equation where two times the value of the paper-ad revenue is less than that of the online ad revenue, then you can justify the leading market executive's claims in 2) part A

unicorn35:

which means everything in my information of showing my work was right but u just told me i didnt need that -_- see now you confusing me

unicorn35:

if you could kindly tell me what im missing that would help alot

Jaynater:

@unicorn35 wrote:
which means everything in my information of showing my work was right but u just told me i didnt need that -_- see now you confusing me
I didn't mean it like that. I suppose we both got a little confused. Yes, you do need your work, but you didn't quite need it for the question I was trying to help you answer. So, yes you're correct in your work. Anyways, you said you need help with 3?

Jaynater:

The bottom part, right? I can help you with that if that's what you need.